Show that the circumference of the ellipse (π‘₯ 2/ π‘Ž2 )+( 𝑦 2/ 𝑏2) = 1Β is given by 2πœ‹π‘Ž [1- βˆ‘βˆžπ‘›=1 ( (2π‘›βˆ’1)!!/ (2𝑛)!! ) 2( 𝑒 2𝑛 /2π‘›βˆ’1 ) ]. Here 𝑒 = √(1 βˆ’( 𝑏2/ π‘Ž2 )), π‘Ž > 𝑏 is the eccentricity. Length of graph 𝑦 = 𝑓(π‘₯) can be found by ∫ √(1 + ( 𝑑𝑦 /𝑑π‘₯) 2 )𝑑x.


1
Expert's answer
2021-11-11T11:30:38-0500

An ellipse equation is given by "\frac{x^2}{a^2}+\frac{y^2}{b^2}=1"

If we put a set of parameters together, we get a parametrization.

"x = a \space cost, y = b \space sin t, 0 ≤ t ≤ 2π."

Calculus's formula for determining the length of a curve is as follows:

"\int_0^{2\pi}\sqrt{(x)^2+(y)^2dt}=\int_0^{2\pi}\sqrt{a^2sin^2t+b^2cos^2tdt}"

We may assume without loss of generality that b β‰₯ a. The expression under the integral can be transformed as

"\sqrt{b^2-(b^2-a^2)sin^2t}=b\sqrt{1-e^2sin^2t}" where "e=\sqrt{b^2-\frac{a^2}{b}}" is called the eccentricity The ellipse's size and shape are described by the parameters b (the length of the bigger semi-axis) and eccentricity.

The length of the ellipse's arc in the first quadrant is sufficient because the ellipse is made up of four such arcs of equal length.

As a result, we must assess the integral.

"\int_0^{\frac{\pi}{2}}\sqrt{1-e^2sin^2tdt}"